OpenStudy (anonymous):

what is 66 1/3% of 39?

OpenStudy (lucaz):

66 1/3 = 199/3 % = (199/3)/100 multiply it by 39

OpenStudy (anonymous):

lucaz, I typed the problem wrong, please redo using 2/3 instead of 1/3 and show the division as well. I'm new at this. If anyone else wants to help please feel free. Thanks

OpenStudy (lucaz):

66 2/3?

OpenStudy (anonymous):

yes

OpenStudy (lucaz):

you have to multiply 66 by 3 and add 2, 66 2/3 = 200/3 (200/3)% of 39 is (200/3)/100 x 39 (200/300) x 39 = 2/3 x 39 = (2x39)/3 = 26

OpenStudy (anonymous):

Yay I had it right!! Thanks so much lucaz!

OpenStudy (lucaz):

=]